# Zeroth-order Informed Fine-Tuning for Diffusion Model: A Recursive Likelihood Ratio Optimizer

### v2

@@ 4.2 Variance Analysis
5 Recursive Likelihood Ratio Optimizer
As summarized by Table 1, in DM fine-tuning, the full BP imposes prohibitively high computational and memory requirements, while RL suffers from excessively high variance. In this section, we propose the ZO-informed RLR optimizer, which integrates the above three gradient estimators within a recursive framework. The RLR estimator is an unbiased gradient estimator that pursues minimal variance under a “feasible” computational and memory budget. The RLR estimator is constituted of three parts and takes the form:
$\underbrace{{\frac{\partial\varphi(\bm{x}_{1};\theta)}{\partial\theta}}^{\top}\frac{dR(\bm{x}_{0})}{d\bm{x}_{0}}}_{\mathrm{One-step\ first-order\ estimator}}$ (11)
$+\underbrace{R(\bm{x}_{0}){D^{\top}_{\theta}\phi_{j:j+h}(\bm{x}_{j+h};\theta)}\nabla\ln f(\bm{z}_{j})}_{h-\mathrm{length\ half-order\ estimator}}$
$+\underbrace{\sum_{i\in C}R(\bm{x}_{0})\nabla\ln f(\bm{z}_{i})}_{\mathrm{zeroth-order\ estimator}},$
where $j\sim\mathcal{U}(1,T-h)=\mathcal{U}(\{1,2,...,T-h\})$, $C=\{1,2,\dots,T\}\setminus\{j,j+1,\dots,j+h\}$, $\bm{z}=\{\bm{z}_{i},\bm{z}_{j}\}$ and $\bm{z}_{i}\sim\mathcal{N}(0,\sigma_{i}\bm{I}),\bm{z}_{j}\sim\mathcal{N}(0,\sigma_{j}\bm{I})$ independently.
@@ 5 Recursive Likelihood Ratio Optimizer
In the first part of the RLR estimator, we differentiate the reward model and the first time-step of the diffusive chain, as shown in Figure 2. With the availability of the reward model structure, it is unnecessary to use zeroth-order methods to bypass BP.
@@
The generation process of the DM follows a coarse-to-fine structure, with each time-step in the chain controlling a different scale of generation. Incorporating precise gradient information from every time-step is essential, but full BP is computationally prohibitive. Truncated BP introduces bias, while zeroth-order optimization leads to high variance by ignoring structural information. To address this, the second part of RLR applies half-order optimization on a randomly selected $h$-length sub-chain, ensuring unbiasedness while minimizing variance. Specifically, before each optimization step, a time-step $j\sim\mathcal{U}(1,T-h)$ is uniformly sampled, natural perturbations are applied to the corresponding latent variable, and localized BP is performed over the $h$-length sub-chain, where the hype-parameter $h$ is chosen based on the available computation budget.
@@
For the remaining times steps, $C=\{1,2,\dots,T\}\setminus\{j,j+1,\dots,j+h\}$, we inject noise directly into the model’s parameters to construct an ZO estimator. This approach does not require caching intermediate latent variables or performing BP, making it computationally cheap.
Overall, the RLR reorganizes the recursive computation chain by perturbation-based estimation, seamlessly integrating zeroth-order, half-order, and first-order optimization techniques. RLR strikes a balance between computational cost and gradient accuracy, achieving both unbiasedness and low variance. The following Theorem 5.1 establishes its unbiasedness.
Theorem 5.1.
The RLR estimator is an unbiased estimator:
$\nabla_{\theta}\mathbb{E}[R(\phi_{1:T}(\bm{x}_{T};\theta))]=\mathbb{E}_{\bm{z}\sim f(\bm{z}),j\sim\mathcal{U}(1,T-h)}$ (12)
$\bigg{[}\frac{\partial\varphi(\bm{x}_{1};\theta)}{\partial\theta}^{\top}\frac{dR(\bm{x}_{0})}{d\bm{x}_{0}}+\sum_{i\in C}R(\bm{x}_{0})\nabla\ln f(\bm{z}_{i})$
$+R(\bm{x}_{0})D^{\top}_{\theta}\phi_{j:j+h}(\bm{x}_{j+h};\theta)\nabla\ln f(\bm{z}_{j})\bigg{]}.$
Theorem 5.2.
@@ 6 Experiments
RLR 97.53 97.19 98.05 70.69 66.15 71.08 84.61
We verify the superiority and applicability of the RLR optimizer against various baselines on two generation tasks: Text2Image and Text2Video. Finally, we propose a novel prompt technique that is natural for the RLR optimizer to further demonstrate the intuition of RLR and applicability.
6.1 Setting
Prompts dataset.
We use Pick-a-Pic (Kirstain et al., 2023) and HPD v2 (Wu et al., 2023) for the Text2Image experiments. We report the performance of the trained model on unseen prompts from the training phase. For the Text2Video task, we prompt ChatGPT to generate a series of prompts that describe motions and train models under the prompts. After the training, we evaluate the model performance on the unseen prompts from the vbench (Huang et al., 2024).
Reward model and benchmark.
We adopt multiple human preference reward models to train and test our methods, including PickScore (Kirstain et al., 2023), HPS v2 (Wu et al., 2023), and ImageReward (Xu et al., 2024b). All the models are trained on large-scale preference datasets. We also included the traditional aesthetic models (Schuhmann, 2022), i.e. AES.
For the video generation task, we use the VBench (Huang et al., 2024) to rate the methods according to various perspectives of the generated videos. The generated videos will be evaluated from 6 aspects: Subject Consistency (SC), Background Consistency (BC), Motion Smoothness (MS), Dynamic Degree (DD), Aesthetic Quality (AQ), and Imaging Quality (IQ).
Baselines.
6.2 Text2Image Generation
@@
6.2.1 Sample Efficiency Analysis
The compare the sample efficiency and the variance of different methods, we show the reward curves of SD 1.4 when training on the AES and HPS v2 models in Figure 4. The DDPO optimizes the reward in a very slow pace, indicating high variance and low sample efficiency. In the earlier phase, the AlignProp has comparable performance as the RLR. In the later phase, while the RLR can continue to improve the reward, the AlignProp suffers from severe model collapse.
@@ 6.4 Diffusive Chain-of-Thought
Furthermore, we propose the Diffusive Chain-of-Thought (DCoT), a prompt technique, that is natural for our RLR optimizer to demonstrate the applicability of our method.
It is worth noting that the generation process of the DM is essentially a multi-scale or coarse-to-fine process, first generating global structures and then focusing on local details. We propose dividing all the diffusion process steps into three groups: coarse-level, mid-level, and fine-level. The coarse-level chain includes steps adjacent to the initial noise, focusing on generating a rough outline. The fine-level chain includes steps adjacent to the final output, focusing on the fine-grained details. The mid-level chain in between focuses on the geometric structure of the content.
The idea of DCoT is shown in Figure 8, which converts the original prompt into multi-scale prompts reflecting the coarse-to-fine nature. Different generation steps should conditioned on different prompts instead of conditioned on the same prompt.
@@
The HO estimator term, having lower variance than the zeroth-order term, in the RLR enables a $h$-step local computational chain. As shown in Formula 11, the HO estimator term uniformly picks a starting point $j\sim\mathcal{U}(1,T-h)$ from the entire $T$-step chain to start the local $h$-step BP chain. When applying the DCoT to the fine-tuning process, we should constrain the sample range of $j$ in the steps where deficiencies exist, $j\sim\mathcal{U}(a,b),1<a<b<T-h,b-a>h$. In our experiment for the hand task, we set $a=30$ and $b=40$.
We write 5 prompts for hand generation and then prompt ChatGPT to generate the multi-scale prompts for the three levels. We report the performance in Table 4. As shown in the table, either simply applying DCoT to the base model or combining it with the RLR can improve the performance significantly. The RLR with the DCoT has the best performance. Qualitative results are shown in the appendix.
@@ 6.5 Ablation Study
We conduct the ablation study, using SD 1.4 and HPD v2, to verify the contribution of different parts in the RLR optimizer. In Table 5, we evaluate the RLR and its variant (V1: the RLR without HO and ZO; V2: the RLR without ZO; V3: the RLR without HO). The V1 performs the worst since it actually reduces to the truncated BP with only one time-step. The V2 and V3 perform better than the V1. It is worth noting that the V2 is better than the V3. The V2 without HO is actually an unbiased estimator since it takes all time steps into account when estimating the gradient. Even though the V2 rearranges the computational graph by the HO, it still a biased estimator. This phenomena indicates the importance of unbiasness when conducting the fine-tuning task.
Table 5: Ablation of the RLR
Methods PickScore HPSv2 AES ImageReward RLR w/o HO and ZO 18.43 23.66 5.78 60.07 RLR w/o ZO 19.28 26.70 5.92 63.85 RLR w/o HO 20.11 27.07 6.23 68.35 RLR 21.38 29.22 6.65 76.55
@@ 7 Conclusion
In this paper, we propose a novel zeroth-informed optimizer for the alignment of the DM. We give a perturbation-based analysis for the diffusion chain and draw theoretical insight into the bias, variance, and convergence of the RLR estimator and optimizer. The empirical study on Text2Image and Text2Video is conducted. Our RLR outperforms all baselines under various metrics. Furthermore, we propose a prompt technique that is natural for the RLR optimizer to validate the applicability.
Appendix A appendix
A.1 Experiments settings
A.1.1 Prompts

### v3

@@ 3.2 $\mathcal{G}$: The Unbiased Gradient Estimator Design Space in Diffusion Models
where each sequence $G\in\mathcal{G}_{\text{full}}$ represents a composite estimator composed of local choices at each time step. Table 1 summarizes the variance, unbiasedness, and memory costs of each strategy: FO has the lowest variance but highest cost, ZO is the cheapest but suffers from the highest variance, and HO balances the two. The theoretical underpinnings of these variances, unbiasedness, and memory comparisons are formalized in Section 6, and constitute an independent contribution of this work.
4 Solving the Variance Minimization Problem: Recursive Likelihood Ratio Optimizer
4.1 Recursive Likelihood Ratio Optimizer
Having characterized the decision space $\mathcal{G}_{\text{full}}$, we now turn to solving the variance minimization problem (4). Due to practical constraints, we further reduce the full design space by enforcing the following structure: (i) All HO estimators should be connected in the chain, since separating the HO path would incur high variance. (ii) FO should be directly connected to the reward model, according to its definition. This way, the decision space $\mathcal{G}_{\text{full}}$ is reduce to $\mathcal{G}_{\text{RLR}}=\{(g_{1}^{\text{FO}},\cdots,g_{j}^{\text{HO}},\cdots,g_{j+h}^{\text{HO}},\cdots,g_{T}^{\text{ZO}})\mid 1\leq j\leq T-h\}$, which consists of one FO at the first step, a HO sub-chain of length $h$ starting at step $j$, and ZO estimators for all remaining steps to ensure unbiasedness. Under this specific structure, the decision variables are reduced to the length $h$ and the starting index $j$. Each solution in $\mathcal{G}_{\text{RLR}}$ is referred to as a Recursive Likelihood Ratio (RLR) estimator (see Figure 4), which takes the form
$G$ $=\underbrace{{D^{\top}_{\theta}\phi(x_{1},z_{1};\theta)}\frac{dR(x_{0})}{dx_{0}}}_{\mathrm{One-step\ first-order\ estimator}}$ (6)
$-\underbrace{R(x_{0}){D^{\top}_{\theta}\phi_{j:j+h}(x_{j+h},z_{j:j+h};\theta)}\nabla_{z}\ln f(z_{j})}_{h-\mathrm{length\ half-order\ estimator}}-\underbrace{\sum_{i\in C}R(x_{0})\nabla_{z}\ln f(z_{i})}_{\mathrm{zeroth-order\ estimator}},$
where $j\in[1,T-h]\cap\mathbb{Z}$ , $C=\{1,2,\dots,T\}\setminus\{j,j+1,\dots,j+h\}$, $z_{j}\sim\mathcal{N}(\bm{0},\sigma_{j}I)$, and $z_{i}\sim\mathcal{N}(\bm{0},\sigma_{i}I)$ for $i\in C$.
@@ 4.1 Recursive Likelihood Ratio Optimizer
In the first part of the RLR estimator, we apply the FO estimator to the first time step to directly backpropagate through the reward model, avoiding black-box treatment (e.g., RL and ZO) and better leveraging its structure, as shown in Figure 4.
@@
The generation process of the DM follows a coarse-to-fine structure, with every time step in the chain controlling a different scale of generation. Incorporating precise gradient information from every time step is essential, but full BP is computationally prohibitive. Truncated BP introduces bias, while ZO and RL lead to high variance by ignoring structural information. To address this, the RLR optimizer incorporates an HO $h$-length sub-chain, capturing multi-scale information while minimizing variance. Specifically, the starting index of HO, $j\sim\mathcal{J}(1,T-h)$, is randomly selected across the whole diffusive chain, following a given distribution $\mathcal{J}$ (see Section 3.2). The inherent perturbation, $z_{j}$, enables the localized $h$-length sub-chain, $D_{\theta}\phi_{t:t+h-1}$, effectively capturing the visual scale information represented around that step (see Section 3.2 for choosing $h$).
@@
For the remaining times steps, $C=\{1,2,\dots,T\}\setminus\{j,j+1,\dots,j+h\}$, we inject noise directly into the model’s parameters to construct ZO estimation to ensure unbiasedness. This approach is computationally cheap without caching intermediate latent variables.
4.2 Optimizing $h$ and $j$: Variance–Memory Tradeoffs
The remaining task is to optimize the two variables in the RLR estimator, $h$ and $j$, to solve the optimization problem (4). Notably, the choice of $j$ does not directly affect this surrogate objective, but instead influences the ability to capture multi-scale information across different steps, so we treat its decision as a separate problem of interest.
Optimizing $h$.
To reduce the number of problem parameters that need to be estimated, we use an upper bound on the variance of the RLR estimator (see the Appendix J.4) as a surrogate objective:
$\min_{h\in\mathbb{N}_{0}:\ G(h)\in\mathcal{G}_{\mathrm{RLR}}}$ ${\sum_{t=1}^{T}\text{Var}(g_{t})}+2\sum_{t\neq t^{\prime}}\sqrt{\mathrm{Var}(g_{t})\mathrm{Var}(g_{t^{\prime}})}$ (7)
s.t. $\quad\mathcal{B}_{h}h+\mathcal{B}_{z}(T-1-h)\leq\mathcal{B},$
where $h$ and (T1h)(T-1-h) are the number of steps for HO and ZO; $\mathcal{B}_{h}$ and $\mathcal{B}_{z}$ are coefficients indicating the magnitude of the memory cost of HO and ZO per step. In practice, the available budget satisfies $\mathcal{B}_{z}T<\mathcal{B}<\mathcal{B}_{h}T$, meaning that using pure ZO underutilizes the budget, while using pure HO exceeds it. Let $V^{2}_{h}$ and $V^{2}_{z}$ denote the per-step variance of the HO and ZO, respectively. Since HO and FO have much lower variance than ZO, we use a common $V_{h}$ for both HO and FO, and assume $V_{h}\ll V_{z}$ and $T>2$. These conditions ensure the optimization problem admits the solution:
$h^{\ast}=\min\{\lfloor\frac{\mathcal{B}-\mathcal{B}_{z}(T-1)}{\mathcal{B}_{h}-\mathcal{B}_{z}}\rfloor,\lfloor\frac{TV_{z}}{2(V_{z}-V_{h})}-1\rfloor\}>0.$ (8)
We set $\mathcal{B}_{h}=8\text{GB}$ and $\mathcal{B}_{z}=0.24\text{GB}$, which is supported by empirical evidence in Table 9 in the Appendix. If the memory budget $\mathcal{B}$ is between $30\text{GB}$ and $40\text{GB}$. It is recommended to set $h=2$. As the formula (7) indicates, the variance decreases as $h$ increases. However, the performance exhibits diminishing improvement with increasing $h$. In other aspects, the memory consumption grows linearly with $h$, and the computational time grows even more rapidly. The above claims are corroborated by our ablation results in Table 9. Therefore, even with a larger memory budget, blindly increasing $h$ is not advisable. Moreover, since $V_{h}\ll V_{z}$, the second term in (8) simplifies to approximately $\frac{T}{2}-1\approx 24$, which is typically larger than the first term. As a result, we have $h^{*}=\left\lfloor{(\mathcal{B}-\mathcal{B}_{z}(T-1))}/{\mathcal{B}_{h}-\mathcal{B}_{z}}\right\rfloor$ in practice, and there is no need to estimate $V_{h}$ and $V_{z}$.
Determining $j$.
We use the gradient norm to represent the importance of different steps. Then sample $j$ from the categorical distribution $j\sim\mathcal{J}(1,T-h)=\mathcal{CAT}(\text{softmax}(\|g_{1}\|,\cdots,\|g_{T-h}\|))$.
@@ 5 Experiments
We verify the superiority and applicability of the RLR optimizer against various baselines on two generation tasks: Text2Image and Text2Video. We compare the RLR with the RL-based method (DDPO), and the truncated-BP-based methods (Alignprop and VADER). Moreover, other baselines, e.g, closed-source models, are also included. Finally, we propose a novel prompt technique that is natural for the RLR optimizer, demonstrating the enhanced capability of the proposed RLR optimizer to capture multi-scale information for visual generation. Ablations are included to verify the validity of the proposed RLR optimizer. Please refer to Appendix B for detailed settings.
@@
RLR 97.64 97.19 98.05 70.69 66.15 71.08 84.63
@@
5.1 Text2Image Generation
Sample efficiency analysis.
The compare the sample efficiency and the variance of different methods, we show the reward curves of SD 1.4 when training on the AES and HPS v2 models in Figure 5. The DDPO optimizes the reward at a very slow pace, indicating high variance and low sample efficiency. In the earlier phase, the AlignProp has comparable performance as the RLR. In the later phase, while the RLR can continue to improve the reward, the AlignProp suffers from severe model collapse.
@@ 5.3 Diffusive Chain-of-Thought
Furthermore, we propose the Diffusive Chain-of-Thought (DCoT), a prompt technique that is natural for our RLR optimizer to demonstrate the applicability of our method. The core idea is that DMs generate content in a multi-scale (coarse-to-fine) manner, and deficiencies at a particular scale can be addressed by focusing gradient updates on the corresponding steps of the diffusion process by the HO sub-chain. We propose dividing all the diffusion process steps into three groups: coarse-level, mid-level, and fine-level. The coarse-level chain includes steps adjacent to the initial noise, focusing on generating a rough outline. The fine-level chain includes steps adjacent to the final output, focusing on the fine-grained details. The mid-level chain in between focuses on the geometric structure of the content.
@@
The idea of DCoT is shown in Figure 9, which converts the original prompt into multi-scale prompts reflecting the coarse-to-fine nature. Different generation steps should be conditioned on different prompts instead of being conditioned on the same prompt. The HO estimator term in the RLR enables a $h$-step local computational chain for low-variance, unbiased gradient estimation. By integrating DCoT, we can target the HO sub-chain precisely at the time steps (i.e., scales) where generation is deficient, as revealed by the multi-scale prompt decomposition. The HO estimator term uniformly picks a starting point $j\sim\mathcal{U}(1,T-h)$ from the entire $T$-step chain to start the local $h$-step BP chain. When applying the DCoT to the fine-tuning process, we should constrain the sample range of $j$ in the steps where deficiencies exist, $j\sim\mathcal{U}(a,b),1<a<b<T-h,b-a>h$. In our experiment for the hand task, we set $a=30$ and $b=40$.
We write 5 prompts for hand generation and then prompt ChatGPT to generate the multi-scale prompts for the three levels. We report the performance in Table 4. As shown in the table, either simply applying DCoT to the base model or combining it with the RLR can improve the performance significantly. The RLR with the DCoT has the best performance. Qualitative results are shown in the Appendix F.
@@ 5.4 Ablation Study
Table 5: Ablation of the RLR.
Methods PickScore HPSv2 AES ImageReward RLR w/o HO & ZO 18.43 23.66 5.78 60.07 RLR w/o ZO 20.11 27.07 6.23 68.35 RLR w/o HO 19.28 26.70 5.92 63.85 RLR 21.38 29.22 6.65 76.55
We conduct the ablation study, using SD 1.4 and HPD v2, to verify the contribution of different parts in the RLR optimizer. In Table 5, we evaluate the RLR and its variants (V1: the RLR without HO and ZO; V2: the RLR without ZO; V3: the RLR without HO). The V1 performs the worst since it actually reduces to the truncated BP with only one time-step. The V2 and V3 perform better than the V1. It is worth noting that the V2 is better than the V3. The V3 without HO is actually an unbiased estimator since it takes all time steps into account when estimating the gradient. Even though the V2 rearranges the computational graph by the HO, it is still a biased estimator. This phenomenon indicates the importance of unbiasedness when conducting the fine-tuning task.
6 Theoretical Properties of Gradient Estimators: Bias, Variance, and Convergence
In this part, we analyze the bias of truncated BP and compare the variance of different estimators, backing the claim in Table 1. Thanks to the unbiasedness of the RLR estimator, the convergence of the optimization is also guaranteed.
To alleviate the memory burden of full-step BP, the truncated variant is often employed, backpropagating the gradient with only $T^{\prime}$ steps; $T^{\prime}\ll T$. However, the truncation introduces a structural bias into the gradient estimator. We have the following proposition to justify this structural bias.
Proposition 6.1 (Biasedness of Truncated-BP ).
Assume $R$ and $\phi$ are differentiable almost everywhere, $R$ and $\phi_{t}$ have bounded gradients, then the FO estimator is unbiased. However, the truncated BP estimator $\nabla_{\theta}R(\bm{x}_{0})_{\mathrm{truncated}}$ has a structural bias, which can be specified as below:
$\nabla_{\theta}\mathbb{E}[R(x_{0})]-\mathbb{E}[\nabla_{\theta}R(x_{0})_{\mathrm{truncated}}]=\mathbb{E}_{z_{1:T}}\bigg[\bigg(\sum_{i=T^{\prime}+1}^{T}\frac{\partial\phi_{i}(x_{i},z_{i};\theta)}{\partial\theta}\prod_{j=1}^{i-1}\frac{\partial x_{j-1}}{\partial x_{j}}\bigg)^{\top}\frac{dR(x_{0})}{dx_{0}}\bigg].$
Bias in the estimator can lead to suboptimal updates or even training failure, as the truncated gradient may not follow a true descent direction. This can cause two major issues: model collapse and loss of multi-scale information. In contrast, ZO (Spall, 1992) and HO (Jiang et al., 2024) are unbiased.
The stochastic nature of the DM results in the variance of the estimator. As expected, the variance of the FO estimator is lower than that of the HO and ZO estimators because the differentiation leverages the structural information of the neural network. However, BP introduces significant computational and storage overhead. The following proposition demonstrates that this additional cost is, to some extent, justified, as BP leverages accurate internal structures to reduce estimation variance.
Proposition 6.2 (Variance Comparison).
Under Assumptions (A.1-3) in the Appendix, the variance of FO estimators is less than or equal to ZO estimators, i.e.
$\mathrm{Var}(\nabla_{\theta}R(x_{0}))\leq\mathrm{Var}(R(x_{0})\nabla\ln f(z)).$ (9)
Based on the above proposition, it is straightforward to conclude that the variance of the HO estimator is also less than or equal to that of the ZO estimator, as it is essentially an FO estimator with a perturbation at the start of the sub-chain. The Table 1 presents all the gradient computation methods discussed above.
Overall, the RLR reorganizes the recursive computation chain by perturbation-based estimation, seamlessly integrating ZO, HO, and FO optimization techniques. RLR strikes a balance between computational cost and gradient accuracy, achieving both unbiasedness and low variance. The following Theorem 6.3 establishes its unbiasedness.
Theorem 6.3 (Unbiasedness of RLR).
The RLR estimator is an unbiased estimator:
$\nabla_{\theta}$ $\mathbb{E}[R(\phi_{1:T}(\bm{x}_{T};\theta))]=\mathbb{E}_{z_{1:T},j\sim\mathcal{U}(1,T-h)}\bigg[{D^{\top}_{\theta}\phi_{1}(x_{1},z_{1};\theta)}\frac{dR(x_{0})}{dx_{0}}$ (10)
$-R(x_{0}){D^{\top}_{\theta}\phi_{j:j+h}(x_{j+h},z_{j:j+h};\theta)}\nabla_{z}\ln f(z_{j})-\sum_{i\in C}R(x_{0})\nabla_{z}\ln f(z_{i}).\bigg].$
Theorem 6.4 (Convergence Rate).
@@ 7 Conclusion
We propose the RLR optimizer, a half-order gradient estimation framework designed for efficient fine-tuning of diffusion models. Theoretically, we analyze the bias, variance, and convergence of the RLR estimator and formulate a constrained optimization problem to guide its design. Empirically, RLR consistently outperforms both reinforcement learning and truncated backpropagation methods on Text2Image and Text2Video tasks across multiple human preference reward models and benchmarks. Furthermore, we introduce a novel prompt technique, Diffusive Chain-of-Thought (DCoT), which complements the RLR and further boosts performance. Although determining the appropriate sub-chain length $h$ can be nontrivial in practice, we provide both theoretical justification and empirical ablations to guide practitioners in making informed choices.
Appendix Contents
Appendix A Extended Related Works
Diffusion Probabilistic Models.
Denoising Diffusion Model (Ho et al., 2020; Lu et al., 2022) is one of the strongest models for generation tasks, especially for visual generation (Rombach et al., 2022; Peebles & Xie, 2023; Chen et al., 2024a). Extensive research has been conducted from theoretical and empirical perspectives (Song et al., 2020a; Karras et al., 2022; Song et al., 2020b). It has achieved phenomenal success in multi-modality generation, including image, video, audio, and 3D shapes. The DM is trained on enormous images and videos from the internet (Bain et al., 2021; Wang et al., 2023b; Schuhmann et al., 2022). Empowered by modern architecture (Vaswani, 2017), it has a powerful learning capability for Pixel Space Distribution.
Alignment and Post-training.
After pre-training to learn the distribution of the targeted modality (Achiam et al., 2023; Kaplan et al., 2020), post-training is conducted to align the model toward specific preferences or tune the model to optimize a particular objective. RL has been utilized to align the foundation models toward various objectives (Ziegler et al., 2019; Lambert et al., 2022; Black et al., 2023). DPOK (Fan et al., 2024) studies KL regularization when training a separate DM for each prompt. Supervised learning can also be applied to the post-training phase (Rafailov et al., 2024), either optimizing an equivalent objective (Wallace et al., 2024) or directly differentiating the reward model (Clark et al., 2023; Prabhudesai et al., 2023; 2024). D3PO Yang et al. (2024) utilizes the DPO loss to train the DM. Specialist Diffusion Lu et al. (2023) focuses on sample-efficient, few-shot fine-tuning of large pre-trained diffusion models to enable the generation of new visual styles from as few as 5–10 examples. SPIN (Yuan et al., 2024b) introduces a self-play learning paradigm for diffusion models. For DM, most methods use a neural reward model to align the pre-trained model, and there has been a continual effort to design better reward models (He et al., 2024; Xu et al., 2024a; b).
Forward Learning Methods.
After extensive exploration of model training using forward inferences only (see, e.g., Peng et al., 2022; Hinton, 2022), forward-learning methods (Salimans et al., 2017; Malladi et al., 2023; Chen et al., 2023; Jiang et al., 2024) based on stochastic gradient estimation have recently emerged as a promising alternative to classical BP for large-scale machine learning problems (Zhao et al., 2024; Zhang et al., 2024a). Subsequent research (Ren et al., 2025; Chen et al., 2024b) has further optimized computational and storage overhead from various perspectives, achieving greater efficiency.
Appendix B Experiments settings
B.1 Overall Setting
Prompts dataset.
We use Pick-a-Pic (Kirstain et al., 2023) and HPD v2 (Wu et al., 2023) for the Text2Image experiments. We report the performance of the trained model on unseen prompts from the training phase. For the Text2Video task, we prompt ChatGPT to generate a series of prompts that describe motions and train models under the prompts. After the training, we evaluate the model’s performance on the unseen prompts from the VBench (Huang et al., 2024).
Reward model and benchmark.
We adopt multiple human preference reward models to train and test our methods, including PickScore (Kirstain et al., 2023), HPS v2 (Wu et al., 2023), and ImageReward (Xu et al., 2024b). All the models are trained on large-scale preference datasets. We also included the traditional aesthetic models, e.g., AES (Schuhmann, 2022).
For the video generation task, we use the VBench (Huang et al., 2024) to rate the methods according to various perspectives of the generated videos. We report 6 aspects of the generated videos in the main text: Subject Consistency (SC), Background Consistency (BC), Motion Smoothness (MS), Dynamic Degree (DD), Aesthetic Quality (AQ), and Imaging Quality (IQ). For more results of 16 metrics on VBench, please refer to Table 10 in the appendix.
Baselines.
B.2 Prompts